### v1

# GLDesigner: Leveraging Multi-Modal LLMs as Designer for Enhanced Aesthetic Text Glyph Layouts

Abstract
Text logo design heavily relies on the creativity and expertise of professional designers, in which arranging element layouts is one of the most important procedures. However, few attention has been paid to this specific task which needs to take precise textural details and user constraints into consideration, but only on the broader tasks such as document/poster layout generation. In this paper, we propose a VLM-based framework that generates content-aware text logo layouts by integrating multi-modal inputs with user constraints, supporting a more flexible and stable layout design in real-world applications. We introduce two model techniques to reduce the computation for processing multiple glyph images simultaneously, while does not face performance degradation. To support instruction-tuning of out model, we construct two extensive text logo datasets, which are 5x more larger than the existing public dataset. Except for the geometric annotations (e.g. text masks and character recognition), we also compliment with comprehensive layout descriptions in natural language format, for more effective training to have reasoning ability when dealing with complex layouts and custom user constraints. Experimental studies demonstrate the effectiveness of our proposed model and datasets, when comparing with previous methods in various benchmarks to evaluate geometric aesthetics and human preferences. The code and datasets will be publicly available.
1 Introduction
Automatic layout design in creative AI is gaining attention for boosting efficiency and supporting designers. Traditional rule-based methods often produce plain, unappealing layouts and rely on professional designers for aesthetic results, making the process labor-intensive. Particularly, text logo layout designing is challenging and less explored than graphic designs because it requires careful consideration of font style, size, texture and keyword semantics, to ensure the final design is both visually appealing and collision-free.
Figure 1: Illustrated are examples of text logos from our datasets, each glyph uniquely colored. We aim to enhance previous datasets by introducing more complex, compact layouts (top) and extended text sequences (bottom).
To overcome these challenges, recent approaches leverage generative models such as Generative Adversarial Networks (GANs) [10] and Diffusion Models [35] to synthesize graphic layouts by learning from examples of human-designed data. The previous method [38] utilized GAN-based models to address this issue but remains limited in handling (1) long text sequences, (2) user-defined layout constraints, and (3) data variability and generality, making it less suitable for real-world design applications. Additionally, the method encodes character semantics using pretrained semantic word embedding models [26], which are less expressive compared to contemporary large language models (LLMs) that are trained on trillions of text tokens. Meanwhile, an increasing number of general layout generation methods [9, 36, 28, 41] have begun exploring multi-modal, content-aware conditions (e.g., background images, document contexts). However, these approaches cannot be directly applied to the text logo layout task, as it requires managing multiple glyphs and incorporating more fine-grained details to ensure both aesthetic appeal and readability.
Acknowledging these limitations, we present a unified VLM-based framework for text logo layout generation, leveraging their inherent ability to handle multi-modal inputs of glyph images and textual semantics. We employ pre-trained VLMs [29] with minor modifications, including early feature extraction to enhance sensitivity to fine-grained textural details and prevent collisions between glyphs. To improve computational efficiency, we apply adaptive average pooling, reducing the number of visual glyph tokens by a factor of $6^{2}$, which accelerates both training and inference without adding extra parameters or compromising performance. The final layout outputs are represented as pure JSON texts, ensuring a standardized format that facilitates easy and stable post-rendering while enhancing training stability. Additionally, we fine-tune the LLM with instruction-following techniques for the text logo layout task, enabling it to create aesthetically pleasing layouts by analyzing glyph styles, semantic meanings, and optionally incorporating user constraints.
We construct two large-scale text logo datasets featuring pixel-level annotations, detailed layout descriptions, and user constraints, comprising over 17k samples, which is six times larger than existing public datasets [38]. These datasets include synthesized logos rendered by pre-defined rules from human designers and general text logos sourced from the internet, which are subsequently subjected to post-processing and annotation, as illustrated in Fig. 1. Extensive experiments demonstrate that leveraging both datasets significantly reduces glyph collisions and enhances layout creativity and diversity. Our approach outperforms all existing methods on qualitative benchmarks and human preference evaluations while also satisfying additional layout constraints, enabling more controllable generation in real-world applications.
The main contributions of this work are as follows:
  1. 1.
    We introduce a VLM-based framework that generates content-aware text logo layouts by integrating multi-modal inputs and optionally supports user constraints through natural language instructions, making it well-suited for real-world design scenarios.
  2. 2.
    We construct two extensive text logo datasets, comprising synthesized and general samples respectively, and generate instruction-following data with fine-grained layout descriptions, significantly surpassing previous efforts in both scale and detail.
  3. 3.
    Our method outperforms existing approaches across multiple benchmarks, and effectively handles long text sequences and user-defined layout requirements. By combining with glyph transformation and textural generation, our model showcases its versatility in real-world applications.
2 Related Work
2.1 Automatic Layout Generation
Graphic Layout Generation encompasses neural networks to automatically arrange elements like images, text, and backgrounds into appealing designs. Early models such as LayoutGAN [24] and LayoutVAE [20] focused on document and poster arrangement without specific content considerations, utilizing generative models [10, 22]. Recent approaches with auto-regressive [2, 11, 23] and diffusion models [18, 44, 17] enhance layout realism and diversity but often lack the content specificity needed for real-world applications. In contrast, content-aware layout generation has grown to address more complex design scenarios, especially in commercial poster design with non-empty background images, guided by saliency maps [14, 46]. Conditional models integrate visual and textual encoders to predict layouts based on specific conditions [4, 43], while other methods employ constrained optimization [21] and intermediate representations [19, 27] to adhere to user-specified constraints.
Text Logo Layout Generation is a specialized area within graphic design, focusing exclusively on arranging pixel-level text glyphs, which requires careful consideration of glyph strokes and reading order to ensure both aesthetic appeal and readability. Currently, the only existing method [38] utilizes a GAN-based model to integrate both linguistic and visual information from input texts and glyph images to predict optimal layouts. While demonstrating promising performance, it still struggles with designing layouts for long text sequences, accommodating user-defined constraints, and offering diverse layout designs due to the limited amount of training data.
Figure 2: Pipeline Overview of the Layout Generation Method. For clarity and simplicity, vision-language projection layers and detailed prompt templates are not depicted. Utilizing Early Feature Fusion and Adaptive Average Pooling techniques, as detailed in Section 3.2, ensures precise recognition of the glyph image’s textural details with a reduced number of glyph visual tokens. The JSON-structured output guarantees a consistent and reliable rendering of the final text logo.
2.2 Multi-modal Large Language Models
Multi-modal Large Language Models, especially Vision-Language Models (VLMs), have excelled in numerous visual-language tasks. By leveraging LLMs pretrained on extensive text corpora and using visual features as prompts through instruction tuning, these models achieve impressive performance in applications such as image captioning and Visual Question Answering (VQA).
While layouts, which can be encoded in coding formats such as HTML or JSON, are ideally suited to be processed by pre-trained Large Language Models (LLMs), as domain-specific data can be utilized to strength the code generation ability. Following similar idea, several works use the same method to finetune VLLMs to have better performance on document layout understanding [31, 16]. LayoutNUWA [36] views layout generation as a code generation task and LayoutPrompter [28] leverages the training-free RAG pipeline to strengthen the in-context learning ability. Recently, PosterLLaVA [41] employs structured text and visual instruction tuning to generate layouts under specific visual and textual constraints for poster designs. Meanwhile, some contemporaneous works [6, 5, 9] adopt LLMs as middle-ware for designing and direct layout-to-image generation through diffusion models.
However, these approaches cannot be directly applied to text logo layout generation, because they only consider a single background image and lack of consideration of fine-grained glyph stroke details. We propose GLDesigner to gap this bridge by combining all above advantages, and stimulate MLLM designing ability from fine-tuning on our large-scale constructed datasets.
3 Method
3.1 Preliminaries
A text logo can be represented as a collection of $N$ glyph elements, denoted by $\{(g_{i},b_{i})\}^{N}_{i=0}$. Here, each $g_{i}\in\mathbb{R}^{H_{i}\times W_{i}}$ represents the raw visual component of the glyph image. The corresponding position $b_{i}=(x_{i},y_{i},x_{i}^{\prime},y_{i}^{\prime})$ specifies the glyph’s location using a bounding box defined by its top-left $(x_{i},y_{i})$ and bottom-right $(x_{i}^{\prime},y_{i}^{\prime})$ coordinates. These coordinates are normalized relative to the background’s overall width and height, ensuring consistent placement across different resolutions.
In the task of generating text logo layouts, the inputs include each glyph’s character text $T_{i}$, its corresponding glyph image $g_{i}$, and optionally, user-defined constraints $c$. The objective is to predict the positional bounding box $b_{i}$ for each glyph based on these inputs.
USER:
Can you arrange the given glyph images into a legible and aesthetically pleasing layout?
Make sure there is no unnecessary overlap between the texts. Provide the results in JSON
format, with each text’s position and size represented by a bounding box in the format (left,
top, right, bottom). Each coordinate should be a continuous number between 0 and 1.
The glyph image for each text is "T0": <glyph0>, "T1": <glyph1>, …
The layout requirement is placing all elements in a horizontal line.
Here is the initial JSON: [{"word": "T0", "detail": "", "box": []}, …]
ASSISTANT:
There are N characters in this image, arranged horizontally from left to right ...
[{"word": "T0", "detail": "This character is the largest and positioned on the far left.",
"box": [0.4759, 0.0231, 0.5863, 0.3790]}, …]
Table 1: Prompt templates for applying instruction fine-tuning on the content-aware text logo generation task. The placeholder $T_{i}$ and <glyphi> will be replaced with character texts and glyph visual features during training and inference. Note this is one example of the prompt template, and we have an extensive prompt lists offering unique descriptions of layout-related questions, which will be randomly selected during training to enhance generalization and adaptability.
3.2 Model Architecture
We provide a comprehensive overview of our layout generation pipeline in Fig. 2, with detailed explanations elaborated in the subsequent subsections. The architecture of our model is built upon LLaVA [29], which facilitates the unified processing of glyph images and textual content. Specifically, we utilize ViT-L [33] as the glyph encoder and LLaMA [37] as the LLM. To further enhance the model’s performance, we introduce two innovative techniques: Early Feature Fusion and Adaptive Average Pooling. The first technique enhances the model’s ability to capture and interpret intricate glyph textural details, such as stroke patterns and appearance characteristics. Concurrently, The latter boosts computational efficiency when processing multiple glyph images by reducing the number of glyph visual tokens. These enhancements collectively empower our model to more effectively understand and generate precise layout configurations.
Early Feature Fusion. The default configuration of LLaVA [29] utilizes the visual features from CLIP’s [33] second-to-last layer for vision-language feature projection. While these features focus more on capturing the global properties and semantics of glyph images, but tend to overlook the finer textural details present in the earlier layers. Therefore, we choose to fuse the early visual features into the original features, and experimentally and symmetrically choose the second CLIP layers. We also add a trainable two-layer MLP as the projection layer, which is zero-initialized for training stabilization.
Adaptive Average Pooling. Another limitation of the original model is the linear increase in the number of visual tokens when processing multiple glyph images simultaneously. The number of processed tokens can easily exceed model’s maximum sequence length, leading to decreased training and inference efficiency. To address this issue, we utilize adaptive average pooling to compress the number of visual tokens without introducing additional parameters beyond the existing trainable modules [25, 3]. As demonstrated in [40, 42], this is a parameter-free technique that preserves the spatial locality while retaining essential visual attributes. Moreover, our experimental results confirm that this approach significantly reduces computational overhead during both training and inference phases, all without compromising the quality of layout generation.
3.3 Layout Structure
We introduce tailored prompt templates and a structured textual layout arrangement in JSON format, crafted specifically for the instruction-following fine-tuning stage. Details of the template are provided in Tab. 1.
To effectively guide the model, we begin with a detailed task description in the instruction prompt, which serves to clearly indicate and specify both the task content and the desired output format. For handling glyph images alongside their textual semantics, each glyph is represented with unified tokens as $T_{i}$: <glyphi>, where $T_{i}$ and <glyphi> represent the the $i$-th text character and the visual features of the glyph, as extracted in Sec. 3.2, respectively. Finally, we append an initial empty JSON structure that retains only the essential keys. This minimalist JSON setup helps streamline the training process by providing a clear and uncluttered framework for the model to learn within.
To construct the ground-truth response answers, we enhance the JSON structure by incorporating a comprehensive global layout description alongside a “detail” property. The “detail” property specifies the detailed description for each individual glyph, leveraging the chain-of-thought methodology [39]. This approach facilitates the management of complex and lengthy sequence layouts through a systematic, step-by-step reasoning process. This enhancement not only boosts the accuracy of intricate layout generation but also ensures greater flexibility and responsiveness to a diverse range of user requirements.
4 Datasets
Dataset Train Test Glyphs/img Total Glyphs
TextLogo3K [38] 3170 300 4.69 16274
SynTextLogo 10000 1000 4.36 47963
GenTextLogo 6586 482 5.24 37036
Table 2: Dataset statistics of existing dataset and our datasets. Our proposed datasets boast a significantly larger scale than the existing one, featuring a higher average and total number of glyphs, specifically designed to address complex and lengthy sequence scenarios that [38] struggles to handle.
Figure 3: Textual Layout Description complemented by our data processing pipeline in Sec. 4.3. We provide detailed textual descriptions for the overall text logo (top) and each text glyphs (bottom), with geometric relationships between glyphs.
Given that the existing text logo datasets might be inadequate for effectively fine-tuning VLMs, we embarked on creating two extensive text logo datasets: SynTextLogo and GenTextLogo. SynthTextLogo comprises logos generated from pre-defined templates crafted by skilled designers, ensuring a diverse range of styles and typographies. Meanwhile, GenTextLogo features a curated collection of high-quality, real logos from movie and commercial posters. For each sample, we offer a comprehensive set of annotations, including bounding boxes, pixel-level masks, and character text recognitions, as well as detailed layout descriptions for every character. The subsequent subsections outline our data processing pipeline, complemented by a detailed statistical analysis of the datasets in Tab. 2. We provide examples in Fig. 1 and more can be found in supplementary.
4.1 Synthesized Text Logo Layout
We initially developed a synthesized dataset utilizing text rendering engines, which enabled us to generate diverse styled glyph images from a variety of TTF fonts. Next, we collaborated with human designers to formulate several layout templates, arranging the rendered glyphs according to these templates to craft unique text logos using movie titles for character combinations. To streamline the process, we implemented an automated pipeline that efficiently extracts binary masks and bounding box coordinates for each glyph. While the complete collection comprises over 50k text logos, we selectively sampled 10k for training, ensures the balance between synthesized and general data.
4.2 General Text Logo Layout
To enhance data diversity, aesthetics, and align with real-world design scenarios, we have curated a comprehensive text logo dataset sourced from the internet. This collection features text logos from movie posters and commercial product advertisements, showcasing titles that are intricately stylized and expertly designed, which boasts far more complex layouts compared to traditional text fonts and predefined templates. To streamline the annotation process, we exclusively retain logos with RGBA channels, leveraging the alpha channel to effortlessly derive foreground masks. For each glyph, we meticulously annotate binary masks and text characters, supplemented by the metadata of the logo image. In total, we collect 7,068 logo samples, with 6,586 for training and 482 for testing. In comparison to the previous dataset [38], our collection offers greater diversity across logo types and styles, with longer text sequences that accommodate more complex design scenarios.
4.3 Textual Layout Description
In addition to providing spatial information for each glyph, we enrich our dataset with detailed textual descriptions, which enables a deeper understanding of the overall layout and the geometric relationships between glyphs, as illustrated in Fig. 3. Leveraging off-the-shelf VLMs [1, 32, 3, 7], we automatically generate detailed layout descriptions for each glyph within a text logo, alongside a comprehensive description of the logo’s overall structure and style. Next, we employ a rigorous filtering process to ensure the accuracy of the generated descriptions by cross-verifying them with LLMs [8, 32], through prompting with generated outputs and actual position coordinates, and human annotators, thus retaining only the descriptions that accurately align with the glyphs and correcting those that do not. Most importantly, we ultimately extract and reconstruct user constraints for each logo based on the generated layout descriptions, paving the way for a user-friendly, constraint-conditioned layout generation scenario.
5 Experiments
5.1 Implement Details
Unless otherwise specified, we utilized the pre-trained 7B LLaVA-1.5 model [29] with CLIP-L/336 serving as the glyph visual encoder. Each input glyph image has a size of 336x336, and for these inputs, we implemented two distinct formats: square glyphs that maintain the original aspect ratio by padding to form a square, and tight glyphs that are resized directly. During training, we randomly choose the format type, which aims to enhance the model’s ability to perceive the textural details of glyphs while preserving their original aspect ratio as closely as possible, avoiding excessive distortion. Unlike the approach in [38], which does not impose such constraints, ours seeks a balance between detail accuracy and aspect ratio fidelity. Once the layout results are generated, the entire logo is rendered using the differentiable composition techniques outlined in [38].
To maintain training efficiency while preserving the model’s generalization capabilities, we keep the vision encoder frozen and only fine-tune the LLM and the projection layers. The newly added early feature projection layer is zero-initialized [15, 45] to avoid disruptive noise that could potentially compromise the stability and effectiveness of the pre-trained models.
We conduct training on 8 NVIDIA V100 GPUs, with a total batch size of 128. The AdamW [30] optimizer with a learning rate of 2e-5 and a weight decay of 0.01 is employed, in conjunction with a cosine annealing scheduler that includes 3% warm-up steps. The training process integrates all three datasets for 7 epochs, which completes within 12 hours.
5.2 Evaluation Metric
For a convenient comparison with prior approaches, we utilize the same evaluation metrics as [38] including Fréchet Inception Distance (FID) [13] and Inception Score [34]. These metrics objectively assess the visual quality and diversity of the rendered logos from generated layouts.
However, the aforementioned metric falls short in fully capturing the geometric aesthetics of the generated logos, as it merely assesses the similarity between the generated and original outputs. To provide a more comprehensive evaluation of both geometric and aesthetic qualities, we additionally employ Overlap IoU, Visual Balance [41], and Glyph Ratio Consistency as quantitative metrics.
The Overlap IoU metric evaluates the ratio of pixel-level overlap rather than bounding box-level intersections in document layouts [14, 46]. Visual Balance (V.B) [41] assesses the spatial equilibrium of the generated logo, which ensures that the logo’s positioning is harmoniously centered within the canvas, contributing to an aesthetically pleasing design. Glyph Ratio Consistency (Ratio) calculates the aspect ratio difference between the generated bounding box and the original glyph image, which is an essential metric to keep the original glyph’s shape unchanged.
Most importantly, to assess how well the model adheres to the specified input constraints, we sample a subset of 50 layouts from the test set and ask human annotators to verify if the generated layouts meet the given constraints. Following [41], the average percentage of these violations is referred to the Constraint Violation Ratio (denoted as ViO).
Methods FID $\downarrow$ IS $\uparrow$ IoU $\downarrow$ V.B $\downarrow$ Ratio $\downarrow$
Ground-Truth - - 0.00 4.06 0.00
rule (a) 56.93 1.36 1.28 5.68 0.00
rule (b) 55.46 2.42 2.55 5.45 0.00
[…]
ours 33.48 3.22 17.67 3.78 16.17
$\dagger$ 38.73 3.14 6.21 3.62 22.62
ours (13B) 33.77 3.57 14.12 3.52 14.88
ours (13B)$\dagger$ 38.51 3.14 7.18 3.56 22.88
Table 3: Quantitative results of Unconstrained Layout Generation on the TextLogo3K [38] dataset. $\dagger$ indicates using tightly-cropped glyph images without keeping the aspect ratio, which is the default setting of [38].
Methods FID $\downarrow$ IS $\uparrow$ IoU $\downarrow$ V.B $\downarrow$ Ratio $\downarrow$
Ground-Truth - - 0.00 3.92 0.00
rule (a) 60.23 1.01 10.47 5.21 0.00
rule (b) 56.44 1.38 12.98 5.94 0.00
[…]
ours 26.22 2.38 19.84 3.78 16.72
$\dagger$ 28.76 2.54 8.68 3.61 17.87
ours (13B) 25.23 2.42 20.24 3.80 14.87
ours (13B)$\dagger$ 29.42 2.58 7.86 3.60 17.52
Table 4: Quantitative results of Unconstrained Layout Generation on our GenTextLogo dataset. $\dagger$ indicates using tightly-cropped glyph images without keeping the aspect ratio.
Figure 4: Qualitative comparisons of Unconstrained Layout Generation on the TextLogo3K [38] dataset. $\dagger$ indicates using tightly-cropped glyph images without keeping the aspect ratio. The samples illustrate our model’s enhanced ability to manage long text sequences, surpassing [38].
5.3 Result Comparison
Unconstrained Text Logo Layout Generation. We begin by assessing the capability to produce layouts without constraints across two distinct datasets, depicted in Tab. 3 and Tab. 4. Since our model is able to process two distinct formats of glyph images (as detailed in Sec. 5.1), for testing purposes, we utilize both square-padded and tightly-cropped glyph images, where the latter is marked with $\dagger$ for identification. When evaluating, we remove the layout requirement part from the prompt template, as outlined in Tab. 1, which allows the model full creative freedom in generating layouts.
In our experiments, we also incorporated two rule-based methods, labeled as rule (a) and rule (b) in Tab. 3 and Tab. 4, where glyphs are arranged (a) in a horizontal line and (b) in a horizontal or vertical line randomly. The results illustrate that our model surpasses earlier methods, excelling in both logo quality and geometric aesthetics.
Impressively, our model consistently generates aesthetically pleasing layouts while preserving the original aspect ratio of glyphs, regardless of the input format. This capability offers a significant advantage in processing stylized glyphs, as illustrated in Fig. 4, underscores our model’s perceptual ability to comprehend glyphs across various formats, thereby improving its generalization when handling unseen glyph shapes.
Figure 5: Qualitative comparisons of Unconstrained Layout Generation on our GenTextLogo dataset.
User-Constrained Text Logo Layout Generation. We further conducted experiments to assess the model’s ability to generate outputs based on user-defined constraints, aligning more closely with real-world scenarios. Our findings are presented through both quantitative and qualitative analyses in Tab. 5 (utilizing the ViO metric) and Fig. 6, respectively. Due to the limitations of [38], which does not accommodate input constraints, we executed a prompt-tuning process to ensure a fair comparison. The constraints are encoded using the pre-trained CLIP Text encoder [33], and then integrated into the condition encoder described in [38], serving as an additional condition to manage the constraints effectively.
However, as demonstrated by the ViO metric in Tab. 5, the fine-tuned version of [38] still struggles to comprehend and adhere to user constraints, with a violation rate exceeding 50%, which indicates that the output layout is essentially generated at random. In contrast, our model, which is enhanced with the LLM, exhibits a significantly superior ability to understand and implement user-specified constraints. Moreover, the visualization results in Fig. 6 further highlight our model’s capability to produce diverse layout arrangements while rigorously adhering to the given constraints, showcasing both flexibility and precision.
Figure 6: Qualitative results of User-constrained Layout Generation on our GenTextLogo dataset.
Methods Prefer. $\uparrow$ Quality $\uparrow$ ViO $\downarrow$
TextLogo* [38] 39.5% 2.5 0.57
ours 68.3% 3.4 0.11
ours (13B) 70.9% 3.9 0.09
Table 5: User studies of User-constrained Layout Generation on human annotators.
5.4 Ablation Studies
We conducted extensive ablation experiments to uncover the contributions of each component, as detailed in Tab. 6. By implementing model adjustments such as Early Feature Fusion and Adaptive Average Pooling, we observed notable improvements in mitigating glyph collision reflected in IoU scores, and reduced computational requirements (+2.4% with E.F.F and -86% with A.A.P when processing an average of 1-10 glyphs). Moreover, with the inclusion of additional datasets and textural layout description annotations, we observed significant enhancements in geometric accuracy, aesthetic quality, and layout diversity.
Additionally, augmenting the glyph inputs resulted in improved aspect ratio consistency, effectively preserving the original shapes of highly stylized glyphs, which is crucial for maintaining the integrity of complex designs.. By increasing the model’s scale from 7B to 13B, we observed further performance improvements, offering an optimal balance between computational efficiency and output quality.
Methods FID $\downarrow$ IS $\uparrow$ IoU $\downarrow$ V.B $\downarrow$ Ratio $\downarrow$
baseline 33.86 1.19 29.17 5.97 39.88
+E.F.F 30.46 1.21 27.66 4.63 34.91
+A.A.P 30.77 1.03 26.92 4.66 33.01
+SynTextLogo 28.56 2.04 19.45 3.88 20.81
+GenTextLogo 26.22 2.38 19.84 3.90 16.72
+“detail” 26.00 2.36 19.51 3.78 16.94
Scaling to 13B 25.23 2.42 20.24 3.80 14.87
Table 6: Ablation studies of   model and   training data on GenTextLogo dataset. E.F.F and A.A.P stand for Early Feature Fusion and Adaptive Average Pooling in Sec. 3.2. “detail” stands for adding textural descriptions as illustrated in Tab. 1 and Fig. 3.
Figure 7: Visualization of synthesizing text logos by combing with font generation and texture interpolation [12] in the pipeline.
5.5 User Studies
In addition to evaluating generative capabilities with objective benchmarks, we conducted user studies for a subjective comparison of quality. Mirroring the ViO metric discussed in Sec. 5.2, we asked participants to choose logos they perceived to be of higher quality based on personal preference. The sampled test data included ground-truth logos sourced from the TextLogo3K and GenTextLogo datasets, alongside our generated logos using identical text content but adhering to various layout constraints.
We present the preference rate, defined as the average percentage of instances where participants deemed the generated logos superior to the original ones. Additionally, we provide the average Quality ratings (from 1 to 5, worst to best), as evaluated by participants based on logo quality. The findings, detailed in Tab. 5, demonstrate that our model significantly outperforms previous methods when applied to real-world scenarios, reflecting its enhanced ability to produce visually appealing and user-preferred designs.
We finally showcase a comprehensive pipeline for the automatic generation of text logos, integrating both font generation and textural interpolation [12]. This process begins with the synthesis and rendering of font glyphs based on user-specified text inputs, and prompting our model to generate an optimal layout arrangement. Finally, textural interpolation is applied to enhance the logos with rich semantic textures, elevating their visual appeal. Examples of the pipeline are showcased in Fig. 7, with slight cropping applied for space efficiency.
6 Conclusion
We introduce GLDesigner, a pioneering VLM-based designing framework content-aware text logo layout generation,, capable of creating aesthetic text logos by integrating multi-modal inputs with optional user constraints natural language. We propose two innovative techniques, Early Feature Fusion and Adaptive Average Pooling, designed to significantly reduce the computational load associated with processing multiple glyph images concurrently, which results in faster training and inference without compromising performance. Additionally, we have constructed two comprehensive text logo datasets, SynTextLogo and GenTextLogo, to provide a rich source of training data essential for the instruction fine-tuning phase of our VLM. Our model excels in various benchmarks and user studies, offering a stable, controllable generation of aesthetically pleasing text logos for an enhanced user experience.